A force acts on a 2-kg mass and gives it an acceleration of 3 m/. What acceleration is produced by the same force when acting on a mass of (a) 1 kg? (b) 4 kg? (c) How large is the force?


1
Expert's answer
2022-05-23T10:18:28-0400

The Newton's second law says

"F=ma=\rm 2\: kg*3\: m/s^2=6\: N"

(a)

"a=\frac{F}{m}=\rm\frac{6\: N}{1\: kg}=6\: m/s^2"

(b)

"a=\frac{F}{m}=\rm\frac{6\: N}{4\: kg}=1.5\: m/s^2"